You have been instructed to buy 100 shares in each of the following three initial public offerings.
It has been set a price of € 10 per share for each public offering. The number of shares assigned to you (M which A) and the closing price of the first day is

Share M which A Closing price
A 100 8€
B 80 11€
C 20 15€

What is the total gain or loss on these shares at the close of the first day?
Answer #1

Price of allotment = $10 for each of the three IPO

Gain on Share = Quantity*(ClosingPrice-Allotment Price)

Gain on Share A = 100*(8-10) = -200

Gain on Share B = 80*(11-10) = 80

Gain on Share C = 20*(15-10) = 100

Total Gain = -200+80+100 = -20

So, there is a loss of 20 on the first day